OpenStudy (anonymous):

Factor: 25m^4 - 10m^3 + m^2 +5m - 1

OpenStudy (nikvist):

\[25m^4-10m^3+m^2+5m-1=\]\[=m^2(25m^2-10m+1)+5m-1=m^2(5m-1)^2+5m-1=\]\[=(5m-1)(m^2(5m-1)+1)=(5m-1)(5m^3-m^2+1)\]